OpenStudy (anonymous):

f(x)=4x^2+5x-6; f(x-1)

OpenStudy (anonymous):

Replace every occurrence of x on the right side with x-1 and then simplify. \[f(x-1)=4 x^2-3 x-7 \]

OpenStudy (anonymous):

can u show me

OpenStudy (anonymous):

\[f(x)=4x^2+5x-6\] \[f(x-1)=4(x-1)2+5(x-1)-6

OpenStudy (anonymous):

\[f(x)=4x^2+5x-6\] \[f(x-1)=4(x-1)2+5(x-1)-6

OpenStudy (anonymous):

\[f(x-1)=4(x-1)^2+5(x-1)-6\]

OpenStudy (anonymous):

\[f(x-1)=4(x^2-2x+1)+5x-5-6 \]

OpenStudy (anonymous):

\[f(x-1)=4x^2-8x+4-5x-5-6\]

OpenStudy (anonymous):

f(x-1)=4x^2-3x-7\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

sorry about the gibberish lines also there is a typo where i wrote -5x and it should have been 5x

OpenStudy (anonymous):

welcome